two spherical thread each having a mass of 0.34g and a diameter of 46mm are suspended from a point by threads 40cm long. They are equally charged and repel each other. if the distance between the circumferences of the two charged spherical objects are 24cm. A. calculate the tension on the thread. B. The charge on each spherical object


1
Expert's answer
2021-02-15T17:04:09-0500

a)


"\theta=\arctan{\frac{12}{40}}=16.7\degree"


"T\cos{16.7}=mg\\T\cos{16.7}=(0.34\cdot10^{-3})(9.8)\\T=0.00348\ N"

b)

"8.98\cdot10^9\frac{q^2}{0.24^2}=0.00348\sin{16.7}\\q=8.01\cdot10^{-8}\ C"